OpenStudy (anonymous):

hey i need help solving... 4a-2b+c=-24 a-b+c=13 9a+3b+c=9 can anyone help me with this

OpenStudy (anonymous):

Is this factoring?

OpenStudy (anonymous):

no it wants me to solve it and turn it into a quadratic

OpenStudy (anonymous):

so you need the amount of the variables ?

OpenStudy (anonymous):

yes

OpenStudy (ajprincess):

4a-2b+c=-24-(1) a-b+c=13-(2) 9a+3b+c=9-(3) Let us eliminate c from equations (1) and (2). To do that subtract equation (2) from equation (1). (1)-(2)=> 4a-2b+c-(a-b+c)=24-13 4a-2b+c-a+b-c=11 3a-b=11-(4) Similarly let us eliminate c from equations (2) and (3). To do that subtract equation (2) from equation (3). (3)-(2)=> 9a+3b+c-(a-b+c)=9-13 9a+3b+c-a+b-c=-4 8a+4b=-4-(5) Nw take the equations (4) and (5). 3a-b=11-(4) 8a+4b=-4-(5) Let us eliminate b from both the equations. To do that we need to multiply equation (4) by 4. Then we will get (4)*4=> 4(3a-b)=4*11 12a-4b=44-(6) Nw add both equations (5) and (6) (5)+(6)=> 8a+4b+12a-4b=-4+44 20a=40 Solve for a. Then plug in that value of a in 3a-b=11 to find b. Then plug in the values of a and b in a-b+c=13 to find c. Can u find them @morganalana17?